OpenStudy (anonymous):

PLEASE HELP Part 1: Using complete sentences, explain what the discriminant is and what it tells you about the solutions of a quadratic equation. Part 2: Provide a unique example to back up your explanation.

OpenStudy (anonymous):

1. The discriminant formula is b² - 4ac. It gives information about the nature of the roots of the polynomial. See more in this link: http://en.wikipedia.org/wiki/Discriminan … 2. The example of discriminant that determines the nature of the roots is... x² + 2x + 1 = 0 a = 1 b = 2 c = 1 b² - 4ac 2² - 4*1*1 = 4 - 4 = 0 Since b² - 4ac = 0, this gives us a real root.

OpenStudy (dan815):

quadratic formula comes from completing square method

OpenStudy (anonymous):

so can you help me @dan815